A random sample of

nequals=4949

observations is drawn from a population with a mean equal to

1515

and a standard deviation equal to

1414.

a.

Give the mean and standard deviation of the​ (repeated) sampling distribution

x overbarx.

b.

Describe the shape of the sampling distribution of

x overbarx.

Does your answer depend on the sample​ size?

c.

Calculate the standard normal​ z-score corresponding to a value of

x overbarxequals=12.512.5.

d.

Calculate the standard normal​ z-score corresponding to a value of

x overbarxequals=1919.

a.

mu Subscript x overbarμxequals=nothing

sigma Subscript x overbarσxequals=nothing
Answer #1

SolutionA:

xbar=mu=15

sample standard deviation=s=sigma/sqrt(samplesize)=14/sqrt(49)=14/7=2

Solutionb:

here n=49

follows normal distribution as n>30

ANSWER:normal distribution

Solutionc:

zscore for xbar=12.5

z=xbar-mu/s

=12.5-15/2

=-1.25

ANSWER:

z score=-1.25

Solutiond:

zscore for xbar=19

z=xbar-mu/s

=19-15/2

=4/2

=2

ANSWER:

z score=2